OpenStudy (anonymous):

What is the value of the expression? (11 – 6) • (4 + 3) A. 26 B. 71 C. 12 D. 35

OpenStudy (anonymous):

11-6=5 4+3=7 5*7=35

OpenStudy (anonymous):

Help me plz

OpenStudy (lovelyanna):

11 - 6 = 5 4 + 3 = 7 5 * 7= 35

OpenStudy (lovelyanna):

Which is D

OpenStudy (anonymous):

thank you

OpenStudy (anonymous):

help me with another one?

OpenStudy (anonymous):

What is the value of the expression? (11 – 6) • (4 + 3) A. 26 B. 71 C. 12 D. 35

OpenStudy (anonymous):

but it's the same?

OpenStudy (anonymous):

sorry one sec

OpenStudy (anonymous):

What is the value of the expression when a = 4 and b = 3? 4a + b – 5 A. 7 B. 14 C. 6 D. 2

OpenStudy (anonymous):

4*4+3-5=16-2=14

OpenStudy (anonymous):

Which variable expression represents the word phrase? six decreased by the sum of a number and two A. 6 – (m + 2) B. (m – 2) + 6 C. (m + 2) – 6 D. 6 – (m – 2)

OpenStudy (anonymous):

thank you for helping me :)

OpenStudy (anonymous):

you're welcom i think it's A

OpenStudy (anonymous):

Which equation represents the word sentence? The product of 4 and a number increased by 6 is 30. A. 4x – 6 = 30 B. 4x + 30 = 6 C. 4x + 6 = 30 D. 4x = 30 + 6

OpenStudy (anonymous):

C

OpenStudy (anonymous):

i need help becuse i failed this test before and i don't want to fail again

OpenStudy (anonymous):

i hope you'll pass

OpenStudy (anonymous):

(1)Pt.What is the solution set for the given inequality if the replacement set for r is {5, 6, 7, 8, 9, 10}? 5r ≤ 6r – 8 A. {5, 6, 7} B. {9, 10} C. {8, 9, 10} D. {5, 6, 7, 8}

OpenStudy (anonymous):

What is the solution set for the given inequality if the replacement set for r is {5, 6, 7, 8, 9, 10}? 5r ≤ 6r – 8 A. {5, 6, 7} B. {9, 10} C. {8, 9, 10} D. {5, 6, 7, 8}

OpenStudy (anonymous):